A line with slope equal to 1 and a line with slope equal to 2 intersect at the point (4,6). Points Q and R are both on the x-axis. What is the area of triangle PQR?
Answer by mananth(16949)

m= 1

Plug value of the slope and point ( 4 , 6 ) in
Y = m x + b
6.00 = 4 + b
b= 6 - 4
b= 2
So the equation will be
Y = 1 x + 2
plug y = 0
x=-2
the x intercept will be (0,-2)
--------
m= 2

Plug value of the slope and point (4,6) in
Y = m x + b
6.00 = 8 + b
b= 6 - 8
b= -2
So the equation will be
Y = 2 x -2
for x intercept plug y=0
-2x=-2
x=1
x intercept = (0,1)
You have three points
(0,1), (0,-2),(4,6)
Using the formula find the area of triangle
x1(y3-y2)+x2(y3-y1)+x3(y1-y2)/2
Area = 6 sq. units

m.ananth@hotmail.ca